# Constraint on Neutrino Statistics from Cosmological Data

### v1

@@ Abstract
We investigate the impact of neutrino statistical property on cosmology and the constraints imposed by cosmological data on neutrino statistics. Cosmological data from probes such as Cosmic Microwave Background(CMB) radiation and Baryon Acoustic Oscillation(BAO) are used to constrain the statistical parameter of neutrino. This constraint is closely related to the degeneracy effects among neutrino statistical property, the sum of neutrino masses, and the Hubble constant. Our results show that purely bosonic neutrinos can be ruled out at 95% confidence level and purely fermionic neutrinos are preferred.
@@ 1 Introduction
Neutrinos are among the most elusive and least understood particles in the Standard Model(SM), primarily because they interact very weakly with matter and have extremely small masses. In recent years, significant advancements in precise experiments have greatly improved the constraints on the sum of neutrino masses, $\sum m_{\nu}$. For example, it can be inferred from measurements of neutrino flavor oscillations[1, 2] that
$\begin{array}[]{ll}\sum m_{\nu}\gtrsim 0.06\text{ eV}\quad(\mathrm{NH}),\\ \sum m_{\nu}\gtrsim 0.1\text{ eV}\quad(\mathrm{IH}).\end{array}$ (1)
Cosmology also plays an important role in constraining the sum of neutrino masses. Observations of BAO and CMB are often combined to provide tighter constraints on cosmological models. Since different datasets have different properties, the constraints on $\sum m_{\nu}$ would vary if different combination of datasets is chosen. For instance, the combination of Planck CMB data, WMAP 9-year CMB polarization data and a measurement of BAO from Baryon Oscillation Spectroscopic Survey, Sloan Digital Sky Survey, Wigglez LSS data and the 6dF galaxy redshift survey has provided a constraint $\sum m_{\nu}<0.23\text{ eV}$[3, 4]. When combining the Year 1 (Y1) data from the Dark Energy Spectroscopic Instrument (DESI) with CMB data from Planck PR3 release and the Data Release 6 of Atacama Cosmology Telescope (ACT), an upper limit [5]
@@
is achieved instead. Although this constraint seems more stringent, it may be in tension with limits derived from oscillation experiments.
Besides neutrino mass, the statistical nature of neutrinos is also of great interest. Although neutrinos are generally believed to obey Fermi-Dirac(F-D) statistics, there are very few experimental evidences supporting this hypothesis[6][7]. This makes studying the statistical property of neutrinos from a cosmological perspective important, because modifications to the neutrino statistics can have significant impacts on the evolution of the universe. In some analysis[8, 9, 10, 6] , recent cosmological probes, including the CMB and Big Bang Nucleosynthesis (BBN), have been used to impose constraints on neutrino statistics. However, the bounds are still not stringent enough [8], and the possibility of B-E statistics for neutrinos is allowed at 95% confidence level. Therefore, it is essential to reconsider the statistical property of neutrinos from a cosmological perspective, in particular, to re-analyze the constraints on neutrinos using new data.
@@ 2 Neutrino Statistics and Mass in Cosmology
$\rho_{\nu}=n_{\nu}\int\frac{d^{3}p}{(2\pi)^{3}}f_{\nu}(p)E_{\nu}(p)=n_{\nu}\int_{0}^{\infty}\frac{dp}{2\pi^{2}}\frac{p^{2}E(p)}{\mathrm{exp}(E(p)/T_{\nu})+\kappa_{\nu}}$ (4)
where $n_{\nu}=6$ is the number of degrees of freedom of three flavors of neutrinos and anti-neutrinos. If neutrinos obey F-D or B-E statistics, with $\kappa_{\nu}=1$ or $-1$ respectively, the integral can be calculated analytically using Riemann $\zeta$ function in the relativistic limit,
@@
where $T$ and $\rho_{\gamma}$ are, respectively, the temperature and energy density of photons. In the standard model of cosmic evolution, relativistic neutrinos decouple from the thermal bath earlier than the annihilation of electrons and positrons to photons. The entropy of photons is increased, and, it can be shown that $T_{\nu}/T\approx(4/11)^{1/3}$. For models of variable neutrino statistics, this result is still valid. Therefore, in the relativistic limit, the energy density of B-E neutrino is larger than that of the F-D neutrino by a factor of 8/7. This implies that neutrino statistics can change significantly the total relativistic energy density before they later become non-relativistic in the matter-dominated epoch.
@@
To illustrate the effect of neutrino mass in cosmological evolution, we numerically compute the Hubble expansion rate and give in Fig. 1 the evolution of $H(a)$ for different $\kappa_{\nu}$ where in (a): $\sum m_{\nu}=0$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$ are held fixed; (b): $\sum m_{\nu}=0.06$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$; (c) $\sum m_{\nu}=0.3$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$; (d) $\sum m_{\nu}=0.9$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$. In this figure and the remaining three figures in this section, we fix CMB temperature at present $T_{\mathrm{CMB}}=2.73$ K, baryon energy density $\omega_{\mathrm{b}}\equiv\Omega_{\mathrm{b}}h^{2}=0.0224$ and cold dark matter density $\omega_{\mathrm{cdm}}\equiv\Omega_{\mathrm{cdm}}h^{2}=0.1201$, and we assume the universe is spatially flat. We can see in the figure that smaller $\kappa_{\nu}$ always gives larger $H(a)$ irrespective of the chosen value of $\sum m_{\nu}$ in both radiation-dominated and matter-dominated eras, as expected. This means that smaller $\kappa_{\nu}$ should always give smaller sound horizon, according to Eq.(6). The differences between curves of $\sum m_{\nu}=0$ eV and curves of $\sum m_{\nu}=0.06$ eV are actually very small. One can notice that all the lines approach $1$ as the scale factor $a$ appoaches 1. This is because we have used a fixed $H_{0}$ in these plots which means that the total energy density at the present time has been chosen as a fixed value. One can also notice that the effects of neutrino statistics is more significant in the early radiation-dominated era than in the late matter-dominated era.
@@
Massive neutrinos become non-relativistic in matter-dominated era, contributing to matter energy density $\Omega_{m}$. This results in a significant increase in $H(a)$ in late times. In Fig. 2 , we illustrate the effect of neutrino mass on the evolution of Hubble expansion rate $H(a)$ with (a) $\kappa_{\nu}=1$ and (b) $\kappa_{\nu}=-1$. One can find that a larger $\sum m_{\nu}$ leads to a larger $H(a)$ in the matter-dominated era and in particular near the epoch of recombination. This means that a larger neutrino mass should also lead to a smaller sound horizon, according to Eq.(6). Notice that this effect of neutrino mass is more significant in the matter-dominated era than in the radiation-dominated era. Since the contribution to the sound horizon is determined by the contribution before recombination, the effect of neutrino statistics on the $r_{s}$ is more significant than the effect of neutrino mass. Furthermore, if $\sum m_{\nu}$ is significantly small, for example $\sum m_{\nu}\lesssim 0.1$ eV, neutrino mass would have a very mild effect on the evolution of the universe as shown by the curves for $\sum m_{\nu}=0.1$ eV in Fig. 2.
@@
The above mentioned effects on the sound horizon are illustrated in Fig. 3(a) . The colored and labeled lines are plots for $r_{s}$ versus $\kappa_{\nu}$ with $H_{0}$ fixed to 67 $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$. One can see that smaller $\kappa_{\nu}$ always gives smaller sound horizon irrespective of the chosen value of $\sum m_{\nu}$. The effect of neutrino mass can also be seen clearly in the solid and unlabelled line in Fig. 3(a), for which $\kappa_{\nu}$ is fixed to 1 and $\sum m_{\nu}$ varies from 0 eV to 0.9 eV. We can see that larger neutrino mass gives smaller $r_{s}$, in agreement with the expectation discussed above. But this effect is not as significant as varying the statistics of neutrino, as can be seen in the figure, which is also in agreement with the discussions above.
@@
Figure 4: Evolution of Hubble expansion rate $H(a)$ with the scale factor $a$ for varying the present Hubble expansion rate $H_{0}$. All curves are normalized to the corresponding one with $H_{0}=64$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$. $\kappa_{\nu}=1$ and $\sum m_{\nu}=0.06$ eV are held fixed.
@@ 3 Tools, Datasets and Models
In this section, we briefly describe the prescription of our study, and the datasets and the codes used in our study.
Our study is based on the $\Lambda$CDM model extended by two parameters: neutrino statistics parameter $\kappa_{\nu}$ defined above and the sum of the three degenerate neutrinos $\sum m_{\nu}$. Initially, we fit the model using CMB data and CMB+BAO data respectively to assess whether the latest datasets can impose more stringent constraints on $\kappa_{\nu}$. We refer to the fitting using CMB data as CMB and the fitting using CMB+BAO data as CMB+BAO in later discussions. The details for their data combinations are provided below. To further investigate the degeneracy between $H_{0}$ and $\kappa_{\nu}$, we fit the model with CMB+BAO+$H_{0}$ dataset in which the local measurement on the Hubble rate at present $H_{0}$ is further introduced. We refer to the fitting using this dataset as CMB+BAO+$H_{0}$. This fitting is not ideal enough because of the ’Hubble tension’.
@@
For the CMB data, we combine the baseline temperature (TT) and polarization (EE) auto-spectra, along with their cross-spectra (TE), as incorporated in the Commander likelihood for $\ell<30$ and the plik likelihood for $\ell>30$ from the PR3 release[14]. We further combine the CMB lensing data, which consists of the NPIPE PR4 Planck CMB lensing reconstruction and Data Release 6 from the Atacama Cosmology Telescope (ACT)[15, 16, 17]. The mean value of $H_{0}$ given by Planck PR3 isn’t significantly higher than the one given by Planck PR4. But the constraint has become tighter.
@@
To further study the effects of degeneracy between $H_{0}$ and $\kappa_{\nu}$, local measurement on $H_{0}$ is included for CMB+BAO+$H_{0}$. For a better precision, Pantheon+ analysis of Type Ia supernovae is also imposed to constrain cosmological parameters[21, 22]. It’s not an ideal dataset for precise analysis due to the so-called Hubble tension. However, it could still provide valuable insights for interpreting other results.
The theoretical models are computed with a modification of the Boltzmann code CLASS[13]. All Bayesian inference is performed by Cobaya[23, 24], using Metropolis-Hastings MCMC sampler[25, 26]. For each dataset and model combination, we take advantage of oversampling by setting the parameter oversample_power to 0.4, using the dragging method[27] and running four chains in parallel. If Gelman-Rubin criterion [28] $R-1<0.01$ is satisfied, chains will be stopped. For plotting the posteriors, we use getdist[29].

### v2

@@ Abstract
We investigate the impact of neutrino statistical property on cosmology and the constraints imposed by cosmological data on neutrino statistics. Cosmological data from probes such as Cosmic Microwave Background(CMB) radiation and Baryon Acoustic Oscillation(BAO) are used to constrain the statistical parameter of neutrino. This constraint is closely related to the degeneracy effects among neutrino statistical property, the sum of neutrino masses, and the Hubble constant. Our results show that purely bosonic neutrinos can be ruled out at 95% confidence level.
@@ 1 Introduction
Neutrinos are among the most elusive and least understood particles in the Standard Model(SM), primarily because they interact very weakly with matter and have extremely small masses. In recent years, precision experiments have improved constraints on the neutrino mass differences and hence limit the possible values for their sum. For example, solar neutrino oscillation experiments can measure $\Delta m_{21}^{2}=m_{2}^{2}-m_{1}^{2}$, the mass squared difference of neutrinos in mass eigenstates $\nu_{1}$ and $\nu_{2}$, and atmospheric neutrino oscillation experiments can measure $|\Delta m_{32}^{2}|=|m_{3}^{2}-m_{2}^{2}|$, the absolute value of the mass squared difference of neutrinos in mass eigenstates $\nu_{2}$ and $\nu_{3}$. It can be inferred from these measurements that [1, 2]
$\begin{array}[]{ll}\sum m_{\nu}\gtrsim 0.06\text{ eV}\quad(\mathrm{NH}),\\ \sum m_{\nu}\gtrsim 0.1\text{ eV}\quad(\mathrm{IH}),\end{array}$ (1)
where NH refers to the case with $\Delta m_{32}^{2}>0$ and IH refers to the case with $\Delta m_{32}^{2}<0$. Cosmology also plays an important role in constraining the sum of neutrino masses. Observations of BAO and CMB are often combined to provide tighter constraints on cosmological models. Since different datasets have different properties, the constraints on $\sum m_{\nu}$ would vary if different combination of datasets is chosen. For instance, the combination of Planck CMB data, WMAP 9-year CMB polarization data and a measurement of BAO from Baryon Oscillation Spectroscopic Survey, Sloan Digital Sky Survey, Wigglez LSS data and the 6dF galaxy redshift survey has provided a constraint $\sum m_{\nu}<0.23\text{ eV}$ [3, 4]. When combining the Year 1 (Y1) data from the Dark Energy Spectroscopic Instrument (DESI) with CMB data from Planck PR3 release and the Data Release 6 of Atacama Cosmology Telescope (ACT), an upper limit [5]
@@
is achieved instead. Although this constraint seems more stringent, it may be in tension with limits derived from oscillation experiments as shown in Eq.1.
Besides neutrino mass, the statistical nature of neutrinos is also of great interest. Although neutrinos are generally believed to obey Fermi-Dirac(F-D) statistics, there are very few experimental evidences supporting this hypothesis [6, 7]. This makes studying the statistical property of neutrinos from a cosmological perspective important, because modifications to the neutrino statistics can have significant impacts on the evolution of the universe. In some analysis [8, 9, 10, 6] , recent cosmological probes, including the CMB and Big Bang Nucleosynthesis (BBN), have been used to impose constraints on neutrino statistics. However, the bounds are still not stringent enough [8], and the possibility of B-E statistics for neutrinos is allowed at 95% confidence level. Therefore, it is essential to reconsider the statistical property of neutrinos from a cosmological perspective, in particular, to re-analyze the constraints on neutrinos using new data.
@@ 2 Neutrino Statistics and Mass in Cosmology
$\rho_{\nu}=g_{\nu}\int\frac{d^{3}p}{(2\pi)^{3}}f_{\nu}(p)E_{\nu}(p)=g_{\nu}\int_{0}^{\infty}\frac{dp}{2\pi^{2}}\frac{p^{2}E(p)}{\mathrm{exp}(E(p)/T_{\nu})+\kappa_{\nu}}$ (4)
where $g_{\nu}=6$ is the number of degrees of freedom of three flavors of neutrinos and anti-neutrinos. If neutrinos obey F-D or B-E statistics, with $\kappa_{\nu}=1$ or $-1$ respectively, the integral can be calculated analytically using Riemann $\zeta$ function in the relativistic limit,
@@
where $T$ and $\rho_{\gamma}$ are, respectively, the temperature and energy density of photons. In the standard model of cosmic evolution, relativistic neutrinos decouple from the thermal bath earlier than the annihilation of electrons and positrons to photons. The entropy of photons is increased, and, it can be shown that $T_{\nu}/T\approx(4/11)^{1/3}$. For models of variable neutrino statistics, this result is still valid. Therefore, in the relativistic limit, the energy density of B-E neutrino is larger than that of the F-D neutrino by a factor of 8/7.
In non-relativitic regime, $\rho_{\nu}$ can be approximately written as $\rho_{\nu}\approx m\times n_{\nu}$, where $m$ is the mass of neutrino and $n_{\nu}$ is the number density. The number density at scale factor $a$ is related to the number density at decoupling $n_{\nu}(a_{\mathrm{dec}})$ by
$n_{\nu}(a)a^{3}=n_{\nu}(a_{\mathrm{dec}})a_{\mathrm{dec}}^{3}.$ (6)
Equipped with this we can calculate
$\rho_{\nu}(a)=m_{\nu}\frac{a_{\mathrm{dec}}^{3}}{a^{3}}n_{\nu}(a_{\mathrm{dec}}),$ (7)
where
$\begin{array}[]{ll}n_{\nu}(a_{\mathrm{dec}})&=\displaystyle g_{\nu}\int_{0}^{\infty}\frac{dp}{2\pi^{2}}\,\frac{p}{\exp\bigl{(}E(p)/T_{\nu\mathrm{dec}}\bigr{)}+\kappa_{\nu}}\\ &=\displaystyle\left\{\begin{array}[]{ll}\displaystyle g_{\nu}\,T_{\nu\mathrm{dec}}^{3}/(2\pi^{2})\times\zeta(3)\times\Gamma(3)\times\frac{3}{4},&\text{for }\kappa_{\nu}=1,\\[6.0pt] \displaystyle g_{\nu}\,T_{\nu\mathrm{dec}}^{3}/(2\pi^{2})\times\zeta(3)\times\Gamma(3),&\text{for }\kappa_{\nu}=-1.\end{array}\right.\end{array}.$ (8)
Therefore, in the non-relativistic limit, the energy density of B-E neutrino is larger than that of the F-D neutrino by a factor of 4/3.
Taken together, these factors (8/7 when relativistic and 4/3 when non-relativistic) show how neutrino statistics significantly affect their energy density during the evolution.
@@
To illustrate the effect of neutrino mass in cosmological evolution, we numerically compute the Hubble expansion rate and give in Fig. 1 the evolution of $H(a)$ for different $\kappa_{\nu}$ where in (a): $\sum m_{\nu}=0$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$ are held fixed; (b): $\sum m_{\nu}=0.06$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$; (c) $\sum m_{\nu}=0.3$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$; (d) $\sum m_{\nu}=0.9$ eV and $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$. In this figure and the remaining three figures in this section, we fix CMB temperature at present $T_{\mathrm{CMB}}=2.73$ K, baryon energy density $\omega_{\mathrm{b}}\equiv\Omega_{\mathrm{b}}h^{2}=0.0224$ and cold dark matter density $\omega_{\mathrm{cdm}}\equiv\Omega_{\mathrm{cdm}}h^{2}=0.1201$, and we assume the universe is spatially flat. We can see in the figure that smaller $\kappa_{\nu}$ always gives larger $H(a)$ irrespective of the chosen value of $\sum m_{\nu}$ in both radiation-dominated and matter-dominated eras, as expected. This means that smaller $\kappa_{\nu}$ should always give smaller sound horizon, according to Eq.(9). The differences between curves of $\sum m_{\nu}=0$ eV and curves of $\sum m_{\nu}=0.06$ eV are actually very small. One can notice that all the lines approach $1$ as the scale factor $a$ appoaches 1. This is because we have used a fixed $H_{0}$ in these plots which means that the total energy density at the present time has been chosen as a fixed value. One can also notice that the effects of neutrino statistics is more significant in the early radiation-dominated era than in the late matter-dominated era. Although the variation remains at the level of a few percent even in the radiation-dominated era, it may have important implications in high-precision cosmological data analyses.
@@
Massive neutrinos become non-relativistic in matter-dominated era, contributing to matter energy density $\Omega_{m}$. This results in a significant increase in $H(a)$ in late times. In Fig. 2, we illustrate the effect of neutrino mass on the evolution of Hubble expansion rate $H(a)$ with (a) $\kappa_{\nu}=1$ and (b) $\kappa_{\nu}=-1$. One can find that a larger $\sum m_{\nu}$ leads to a larger $H(a)$ in the matter-dominated era and in particular near the epoch of recombination. This means that a larger neutrino mass should also lead to a smaller sound horizon, according to Eq.(9). Notice that this effect of neutrino mass is more significant in the matter-dominated era than in the radiation-dominated era. Since the contribution to the sound horizon is determined by the contribution before recombination, the effect of neutrino statistics on the $r_{s}$ is more significant than the effect of neutrino mass. Furthermore, if $\sum m_{\nu}$ is significantly small, for example $\sum m_{\nu}\lesssim 0.1$ eV, neutrino mass would have a very mild effect on the evolution of the universe as shown by the curves for $\sum m_{\nu}=0.1$ eV in Fig. 2. This effect is roughly at the level of 0.1%, which is much smaller than the effect induced by $\kappa_{\nu}$ during the radiation-dominated era in Fig. 1.
@@
The above mentioned effects on the sound horizon are illustrated in Fig. 3(a). The colored and labeled lines are plots for $r_{s}$ versus $\kappa_{\nu}$ with $H_{0}$ fixed to 67 $\mathrm{km}$ $\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$. One can see that smaller $\kappa_{\nu}$ always gives smaller sound horizon irrespective of the chosen value of $\sum m_{\nu}$. The effect of neutrino mass can also be seen clearly in the solid and unlabelled line in Fig. 3(a), for which $\kappa_{\nu}$ is fixed to 1 and $\sum m_{\nu}$ varies from 0 eV to 0.9 eV. We can see that larger neutrino mass gives smaller $r_{s}$, in agreement with the expectation discussed above. However, this effect is not as significant as varying the statistics of neutrino, as can be seen in the figure, which is also in agreement with the discussions above.
@@
Figure 4: Evolution of Hubble expansion rate $H(a)$ with the scale factor $a$ for varying the present Hubble expansion rate $H_{0}$. All curves are normalized to the corresponding one with $H_{0}=67$ $\mathrm{km}$ $\mathrm{s}^{-1}\leavevmode\nobreak\ \mathrm{Mpc}^{-1}$, which falls within the $1\sigma$ range of the Planck best-fit estimate, $H_{0}=(67.4\pm 0.5)$ $\mathrm{km}$ $\mathrm{s}^{-1}\leavevmode\nobreak\ \mathrm{Mpc}^{-1}$ [15]. $\kappa_{\nu}=1$ and $\sum m_{\nu}=0.06$ eV are held fixed.
@@ 3 Tools, Datasets and Models
In this section, we briefly describe our analysis prescription, along with the datasets and codes used.
Our study is based on the $\Lambda$CDM model extended by two parameters: neutrino statistics parameter $\kappa_{\nu}$ defined above and degenerate neutrino masses $\sum m_{\nu}$. Initially, we fit the model using CMB data and CMB+BAO data respectively to assess whether the latest datasets can impose more stringent constraints on $\kappa_{\nu}$. We refer to the fitting using CMB data as CMB and the fitting using CMB+BAO data as CMB+BAO in later discussions. The details for their data combinations are provided below. To further investigate the degeneracy between $H_{0}$ and $\kappa_{\nu}$, we fit the model using the CMB+BAO+SNe dataset which includes the Pantheon supernovae sample and a Gaussian constraint on the supernova absolute magnitude from [14]. We refer to the fitting using this dataset as CMB+BAO+SNe. Early-Universe probes, such as CMB typically yield a lower Hubble constant of $H_{0}=(67.4\pm 0.5)$ $\mathrm{km}$ $\mathrm{s}^{-1}$ $\mathrm{Mpc}^{-1}$ [15, 16]. By contrast, local measurements favor a higher value near $H_{0}=(73.0\pm 1.0)$ $\mathrm{km}$ $\mathrm{s}^{-1}$ $\mathrm{Mpc}^{-1}$ [17, 18, 19]. This discrepancy is often referred to as the ’Hubble tension’ [20]. Therefore the resulting constraints should be interpreted with caution, as the combined dataset exhibits significant internal tension.
For the CMB data, we combine the baseline temperature (TT) and polarization (EE) auto-spectra, along with their cross-spectra (TE), as incorporated in the Commander likelihood for $\ell<30$ and the plik likelihood for $\ell>30$ from the PR3 release [15]. We further combine the CMB lensing data, which consists of the NPIPE PR4 Planck CMB lensing reconstruction and Data Release 6 from the Atacama Cosmology Telescope (ACT) [21, 22, 23]. The mean value of $H_{0}$ given by Planck PR3 is not significantly higher than the one given by Planck PR4. However, the constraint has become tighter.
@@
To further study the effects of degeneracy between $H_{0}$ and $\kappa_{\nu}$, local measurement on $H_{0}$ is included for CMB+BAO+Sne. For a better precision, Pantheon analysis of Type Ia supernovae is also imposed to constrain cosmological parameters [14, 27]. Due to the significant Hubble tension, this dataset combination cannot be regarded as statistically well-founded, and its resulting constraints should be interpreted with particular caution. However, it could still provide comparative insights for interpreting other results.
The theoretical models are computed with a modification of the Boltzmann code CLASS [13]. All Bayesian inference is performed by Cobaya [28, 29], using Metropolis-Hastings MCMC sampler [30, 31]. We put flat priors on parameters listed in Table I. For each dataset and model combination, we take advantage of oversampling by setting the parameter oversample_power to 0.4, using the dragging method [32] and running four chains in parallel. If Gelman-Rubin criterion [33] $R-1<0.01$ is satisfied, chains will be stopped. For plotting the posteriors, we use getdist [34].
Parameter Prior range Definition
$\Omega_{\mathrm{b}}h^{2}$ [0.005, 0.1] Baryon density parameter
$\Omega_{\mathrm{c}}h^{2}$ [0.001, 0.99] Cold dark matter density parameter
$100\,\theta_{s}$ [0.5, 10] 100×100$\times$100 × the angular scale of the sound horizon at decoupling
$\tau_{\mathrm{rei}}$ [0.01, 0.8] Optical depth due to reionization
$n_{s}$ [0.8, 1.2] Scalar spectral index
$\ln(10^{10}A_{s})$ [1.61, 3.91] Log power of the scalar power spectrum amplitude
$m_{\nu}$ [0, 1.667] Mass of one of the three degenerate neutrinos
$\kappa_{\nu}$ [-1, 1] Statistical parameter
Table I: The variable parameters and their prior ranges adpoted in the analysis.